# RSCaMa: Remote Sensing Image Change Captioning with State Space Model

### v1

@@ Abstract
Remote Sensing Image Change Captioning (RSICC) aims to identify surface changes in multi-temporal remote sensing images and describe them in natural language. Current methods typically rely on an encoder-decoder architecture and focus on designing a sophisticated neck to process bi-temporal features extracted by the backbone. Recently, State Space Models (SSMs), especially Mamba, have demonstrated outstanding performance in many fields, owing to their efficient feature-selective modelling capability. However, their potential in the RSICC task remains unexplored. In this paper, we introduce Mamba into RSICC and propose a novel approach called RSCaMa (Remote Sensing Change Captioning Mamba). Specifically, we utilize Siamese backbones to extract bi-temporal features, which are then processed through multiple CaMa layers consisting of Spatial Difference-guided SSM (SD-SSM) and Temporal Traveling SSM (TT-SSM). SD-SSM uses differential features to enhance change perception, while TT-SSM promotes bitemporal interactions in a token-wise cross-scanning manner. Experimental results validate the effectiveness of CaMa layers and demonstrate the superior performance of RSCaMa, as well as the potential of Mamba in the RSICC task. Additionally, we systematically compare the effects of three language decoders, including Mamba, GPT-style decoder with causal attention mechanism, and Transformer decoder with cross-attention mechanism. This provides valuable insights for future RSICC research. The code will be available at https://github.com/Chen-Yang-Liu/RSCaMa
@@ I Introduction
Multi-temporal remote sensing images captured by satellites provide rich data resources for surface dynamic monitoring. Remote Sensing Image Change Captioning (RSICC) uses multi-temporal images of the same area to identify and describe surface changes in the form of natural language, such as the categories and attributes of changing objects and the positional relationships between objects. It greatly enhances the readability and practicality of change interpretation information, and has significant application in various fields, including land planning [1], disaster detection [2], and urban expansion research [3].
@@
RSICC involves remote sensing image processing and natural language generation. Current mainstream methods typically employ an encoder-decoder structure. In the encoder, a convolutional neural network (CNN) or Vision Transformer (ViT) is used as the backbone to extract visual features from bi-temporal images, a well-designed “neck” to strengthen the spatial-temporal correlation of bi-temporal features and capture interesting changes features. Finally, the decoder converts these visual features into captions by using language models such as recurrent neural networks or the Transformer.
@@
Most current research focuses on improving the design of the neck module, usually using CNN or Transformer as the basic component, and integrating attention mechanisms to improve the accuracy of change captioning. Hoxha et al. [4] firstly explored the RSICC task and proposed two CNN-based bi-temporal fusion strategies. Liu et al. [5] constructed a large dataset LEVIR-CC and benchmarked several methods such as DUDA [6], MCCFormer [7]. Besides, they proposed a dual-branch Transformer to locate change areas and a Multistage Bitemporal Fusion (MBF) module to perform the fusion of multi-scale features. Chang et al. [8] proposed a hierarchical self-attention module to locate change-related features. PSNet [9] leverages scale-aware reinforcement modules to enhance the perception of changed objects of varying sizes. PromptCC [10] uses an image-level classifier and a feature-level encoder to decouple the RSICC task. Besides, a multi-prompt learning strategy is proposed to exploit a large language model for captioning effectively. CT-Net [11] use employs a Multi-Layer Adaptive Fusion (MAF) module to minimize the impact of irrelevant visual features and a Cross Gated-Attention (CGA) module to attend to multi-scale features during the caption-generation process.
Recently, State Space Models (SSMs) [12] with linear complexity, especially Mamba [13], have demonstrated excellent performance in many language and vision tasks due to their efficient contextual modelling capabilities. In the field of remote sensing, some research has explored the application of SSMs. For instance, Pan-Mamba [14] for pan-sharpening, SpectralMamba [15] for Hyperspectral image classification, MiM-ISTD [16] for infrared small target detection, Samba [17] and RS-Mamba [18] for image segmentation. Two surveys of SSMs can be found in [19, 20].
@@
However, the potential of SSM, particularly Mamba, remains underexplored in the RSICC task. In this paper, we introduce Mamba into the RSICC and propose RSCaMa (Remote Sensing Change Captioning Mamba). Specifically, we utilize the Siamese backbones to extract bi-temporal features, which are subsequently fed into multiple CaMa layers. In each CaMa layer, we propose Spatial Difference-guided SSM (SD-SSM) to enhance spatial change perception, and Temporal Traveling SSM (TT-SSM) to perform temporal interaction between bi-temporal features. The bi-temporal features processed by the CaMa layers will be passed through a projection layer to aggregate bi-temporal information and narrow the gap between the visual and language domains. Finally, language decoders generate sentences describing changes of interest. We experimentally confirm the effectiveness of RSCaMa and demonstrate the potential of Mamba as a new model design component in the RSICC task. Additionally, we compare the effect of three language decoders: Mamba language model, GPT-style decoder with causal attention, and Transformer decoder with cross-attention, providing valuable insight for future RSICC research.
@@
  • We introduce SSM into the RSICC, a multi-modal task, and propose RCaMa for efficient spatial-temporal modelling, providing a benchmark for Mamba-based RSICC.
  • We propose the CaMa layer consisting of SD-SSM and TT-SSM. SD-SSM uses differential features to enhance change perception, while TT-SSM promotes bitemporal interactions in a token-wise cross-scanning manner.
  • The experiment shows the RSCaMa’s SOTA performance and the potential of Mamba in the RSICC task. Besides, we systematically evaluate different language decoding schemes, providing valuable insight for future research.
Figure 1: The illustration of the proposed RSCaMa, which consists of three main components: the backbone, multiple CaMa layers, and the language decoder. The CaMa layers play a pivotal role in significantly enhancing the extraction of discriminative features by integrating SD-SSM and TT-SSM. Specifically, SD-SSM focuses on enhancing change perception in spatial features, while TT-SSM concentrates on temporal interaction.
@@ II-A Preliminaries
$h^{\prime}(t)=\mathbf{A}h(t)+\mathbf{B}x(t)$ (1)
$y(t)=\mathbf{C}h(t)$ (2)
where $h^{\prime}(t)$ refers to the time derivative of $h(t)$, ${A}\in\mathbb{R}^{N\times N}$ is the evolution parameter, ${B}\in\mathbb{R}^{N\times 1}$ and ${C}\in\mathbb{R}^{1\times N}$ serve as the projection parameters.
@@
$\bar{\mathbf{A}}$ $=\mathbf{exp(\Delta A)}$ (3)
$\bar{\mathbf{B}}$ $=\mathbf{(\Delta A)}^{-1}\mathbf{(exp(\mathbf{\Delta}A)-I)\cdot\mathbf{\Delta}B}\approx\mathrm{\mathbf{\Delta}}B$ (4)
@@
$h_{k}=\bar{\mathbf{A}}h_{k-1}+\bar{\mathbf{B}}x_{k}$ (5)
$y_{k}=\mathbf{C}h_{k}$ (6)
Mamba [13] breaks the LTI limitation of SSM by introducing a selective state space mechanism. It allows the matrices $\bar{\mathbf{B}}$, $\bar{\mathbf{C}}$, and $\mathbf{\Delta}$ to vary with the input sequences, enhancing selective information processing across sequences, enabling the model to be contextually aware of the input.
@@ II-B RSCaMa
As depicted in Fig. 1, we present the overall architecture of the proposed RSCaMa. RSCaMa consists of three main components: the backbone, multiple CaMa layers, and the language decoder. Following prior research [10], we employ the CLIP image encoder that has been image-text aligned as the backbone to extract bi-temporal features, which are then added with position embeddings and tokenized as $I_{T_{i}}\in\mathbb{R}^{L\times D}(i=1,2)$. $L$ and $D$ are the number and dimensions of tokens respectively. Subsequently, they are fed into multiple CaMa layers. The CaMa layers leverage the synergistic effect of SD-SSM and TT-SSM to further enhance discriminative feature extraction, with SD-SSM enhancing spatial change perception and TT-SSM concentrating on temporal interaction. The two token sequences are iteratively processed through multiple CaMa layers and then input into a projection layer to further bridge the gap between the visual and language domains. Finally, the language decoder utilizes the resulting embeddings to generate descriptive sentences. For the RSCaMa with $N$ CaMa layers, assuming that the bi-temporal features from the backbone are $I_{T_{1}}^{(0)}$ and $I_{T_{2}}^{(0)}$, the flow of caption generation is expressed as follows:
@@
$\mathrm{cap}$ $=\mathrm{\Phi_{Dec}}(I_{F})$ (9)
where [;] denotes feature concatenation operation on the channel dimension. $\mathrm{\Phi_{Proj}}(\cdot)$ represents a projection layer consisting of a linear layer and a residual convolution block.
@@ II-C SD-SSM
Fig. 2 presents the structure of our proposed SD-SSM. The standard Mamba is designed for one-dimensional sequences. Its causal scanning mechanism makes it unsuitable for spatial-aware understanding in visual tasks. Inspired by [21], we perform Forward-SSM and Backward-SSM scans on flattened visual tokens to achieve bidirectional modelling, enhancing spatial understanding of visual features. Furthermore, to improve the model’s perception of changes, we multiply the bi-temporal differencing features and the output of bidirectional SSMs to guide the model. Let’s denote bi-temporal features input to SD-SSM as $I_{T_{1}}$ and $I_{T_{2}}$, the processing of SD-SSM can be represented as follows:
$I^{o}_{T_{i}}$ $=\mathrm{\Phi_{L}}(I^{F}_{T_{i}})+I_{T_{i}},i=1,2$ (10)
$I^{F}_{T_{i}}$ $=I^{f}_{T_{i}}*I_{d}+I^{b}_{T_{i}}*I_{d}$ (11)
@@
Figure 3: The structure of the TT-SSM, which rearranges two sequences in a bi-temporal token-wise interleaving manner to facilitate temporal interaction.
@@ II-D TT-SSM
SD-SSM spatially interacts with bi-temporal features and differencing features to promote change perception. Furthermore, we propose TT-SSM to perform temporal interaction to further enhance the bi-temporal features. Fig. 3 shows the structure of the TT-SSM. TT-SSM achieves bi-temporal feature interaction by scanning the two token sequences in the time dimension. Specifically, for the two sequences $I_{T_{1}}\in\mathbb{R}^{L\times D}$ and $I_{T_{2}}\in\mathbb{R}^{L\times D}$ from SD-SSM, we rearrange them in a bi-temporal token-wise interleaving manner to construct sequences $I_{s}=[I_{T_{1}}^{(0)},I_{T_{2}}^{(0)},I_{T_{1}}^{(1)},I_{T_{2}}^{(1)},I_{T_{1}}^{(2)},...]\in\mathbb{R}^{2L\times D}$ that effectively facilitate temporal information exchange. Finally, $I_{s}$ will be processed by bidirectional SSM and subsequently split into bi-temporal sequences in the original order. After that, the two sequences will be sent to the next CaMa layer and processed iteratively by SD-SSM and TT-SSM.
@@ III-A1 Dataset
LEVIR-CC [5] is a large-scale RSICC dataset comprising 10,077 pairs of images. These images capture changes over time in 20 distinct regions across Texas, USA, with a temporal range spanning from 5 to 15 years. Each image pair of the dataset has five annotated captions, totaling 50,385 descriptive captions, describing diverse ground surface transformations like building, road, and vegetation.
@@ III-A2 Evaluation Metrics
We employ four quantitative metrics to measure the captioning accuracy, including BLEU-N (N=1,2,3,4), ROUGEL, METEOR, and CIDEr-D. Higher metric scores indicate better model performance. Besides, following [10], we adopt an average metric $S^{*}_{m}$ [22] as follows:
@@ III-B2 Implementation Details
We implemented all models using the PyTorch deep learning framework and trained them on the NVIDIA GTX 4090 GPU. During training, we use the Adam optimizer to optimize the model parameters. The dimension of the word vector is 768. The number of CaMa layers is set to 3. Our final RSCaMa uses one Transformer’s decoder layer with the cross-attention as the language decoder because of its better performance.
@@ III-C Comparison to State-of-the-Art
Currently, mainstream methods commonly utilize CNNs and Transformers as basic deep learning components of the “neck” design to efficiently process bi-temporal features extracted by the backbone. Methods like Capt-Rep-Diff [6], Capt-Att [6], Capt-Dual-Att [6], and DUDA [6] combine CNNs and attention mechanisms to process bi-temporal features and employ LSTM for language decoding. MCCFormer-S [7], MCCFormer-D [7], RSICCFormer [5], PSNet [9] and PromptCC [10] use Transformer as the core component to construct the neck and employ Transformer’s decoder to generate captions. In this section, we compare RSCaMa with these state-of-the-art (SOTA) methods.
As shown in Table I, RSCaMa demonstrates outstanding performance, particularly excelling in key metrics such as BLEU-4 and $S^{*}_{m}$. BLEU-4 measures the similarity of 4-grams between predicted and reference sentences. Specifically, compared to the latest PromptCC, our model shows +1.70% on BLEU-4 and +1.11% on $S^{*}_{m}$. This significant improvement enhancement is attributed to our innovative spatial-temporal modelling units: SD-SSM and TT-SSM. Their collaborative effect enhances the model’s capability to capture spatial-temporal dynamics. Experimental results also suggest that Mamba, as an emerging deep-learning model, has the potential to become a significant option in future model designs. Further exploration of combinations of Mamba with CNNs, Transformers, and Multilayer Perceptrons (MLPs) will be a promising research direction to advance RSICC technology.
TABLE I: Comparisons experiments on the LEVIR-CC dataset, where the bolded results are the best. Different from previous methods relying on CNNs or Transformers as basic elements of the model, combined with various attention mechanisms, our RSCaMa introduce mamba into RSICC for the first time.
Method BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGEL CIDEr-D $S^{*}_{m}$
Capt-Rep-Diff [6] 72.90 61.98 53.62 47.41 34.47 65.64 110.57 64.52
Capt-Att [6] 77.64 67.40 59.24 53.15 36.58 69.73 121.22 70.17
Capt-Dual-Att [6] 79.51 70.57 63.23 57.46 36.56 70.69 124.42 72.28
DUDA [6] 81.44 72.22 64.24 57.79 37.15 71.04 124.32 72.58
MCCFormer-S [7] 79.90 70.26 62.68 56.68 36.17 69.46 120.39 70.68
MCCFormer-D [7] 80.42 70.87 62.86 56.38 37.29 70.32 124.44 72.11
RSICCFormer [5] 84.72 76.27 68.87 62.77 39.61 74.12 134.12 77.65
PSNet [9] 83.86 75.13 67.89 62.11 38.80 73.60 132.62 76.78
PromptCC [10] 83.66 75.73 69.10 63.54 38.82 73.72 136.44 78.13
RSCaMa 85.79 77.99 71.04 65.24 39.91 75.24 136.56 79.24
TABLE II: Ablation studies on the SD-SSM and TT-SSM, where the bolded results are the best.
Method SD-SSM TT-SSM BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGEL CIDEr-D $S^{*}_{m}$
@@
TABLE III: Ablation studies on three language decoders. Our RSCaMa performs best when using the Transformer decoder with cross-attention.

### v3

@@ Abstract
Remote Sensing Image Change Captioning (RSICC) aims to describe surface changes between multi-temporal remote sensing images in language, including the changed object categories, locations, and dynamics of changing objects (e.g., added or disappeared). This poses challenges to spatial and temporal modeling of bi-temporal features. Despite previous methods progressing in the spatial change perception, there are still weaknesses in joint spatial-temporal modeling. To address this, in this paper, we propose a novel RSCaMa model, which achieves efficient joint spatial-temporal modeling through multiple CaMa layers, enabling iterative refinement of bi-temporal features. To achieve efficient spatial modeling, we introduce the recently popular Mamba (a state space model) with a global receptive field and linear complexity into the RSICC task and propose the Spatial Difference-aware SSM (SD-SSM), overcoming limitations of previous CNN- and Transformer-based methods in the receptive field and computational complexity. SD-SSM enhances the model’s ability to capture spatial changes sharply. In terms of efficient temporal modeling, considering the potential correlation between the temporal scanning characteristics of Mamba and the temporality of the RSICC, we propose the Temporal-Traversing SSM (TT-SSM), which scans bi-temporal features in a temporal cross-wise manner, enhancing the model’s temporal understanding and information interaction. Experiments validate the effectiveness of the efficient joint spatial-temporal modeling and demonstrate the outstanding performance of RSCaMa and the potential of the Mamba in the RSICC task. Additionally, we systematically compare three different language decoders, including Mamba, GPT-style decoder, and Transformer decoder, providing valuable insights for future RSICC research. The code will be available at https://github.com/Chen-Yang-Liu/RSCaMa
@@ I Introduction
Multi-temporal remote sensing images captured by satellites provide rich data resources for surface dynamic monitoring [1, 2]. Remote Sensing Image Change Captioning (RSICC) uses multi-temporal images of the same area to identify and describe surface changes in language, such as the changed object categories, the positional relationships between objects, and dynamics of changing objects (e.g., added or disappeared). It greatly enhances the readability of change interpretation information and has significant application in various fields, including land planning, disaster detection, and urban expansion research [3].
@@
RSICC involves remote sensing image processing and natural language generation. Current mainstream methods typically employ an encoder-decoder structure. In the encoder, the CNN or Vision Transformer (ViT) is used as the backbone to extract visual features from bi-temporal images, a well-designed “neck” to enhance bi-temporal feature modeling and capture interesting changes features. Finally, language models are used as the decoder to convert these visual features into captions. Most current research focuses on improving the neck, usually using CNN or Transformer as the basic component, and integrating attention mechanisms to facilitate the accurate perception of changes.
@@
Hoxha et al. [4] firstly explored the RSICC task and proposed two CNN-based bi-temporal fusion strategies. Liu et al. [5] constructed a large dataset LEVIR-CC and benchmarked several methods such as DUDA [6], MCCFormer [7]. Besides, they proposed a dual-branch Transformer to locate change areas and a Multistage Bitemporal Fusion (MBF) module to perform the fusion of multi-scale features. Chang et al. [8] proposed a hierarchical self-attention module to locate change-related features. PSNet [9] leverages scale-aware reinforcement modules to enhance the perception of changed objects of varying sizes. PromptCC [10] uses an image-level classifier and a feature-level encoder to decouple the RSICC task. Besides, a multi-prompt learning strategy is proposed to exploit a large language model for captioning effectively. CT-Net [11] use employs a Multi-Layer Adaptive Fusion (MAF) module to minimize the impact of irrelevant visual features and a Cross Gated-Attention (CGA) module to attend to multi-scale features during the caption-generation process.
Although progress has been made, previous methods have primarily focused on enhancing spatial change perception capabilities, with weaknesses in joint spatial-temporal modeling. Besides, they are constrained by CNNs with limited receptive fields and Transformers with high complexity. Recently, State Space Models (SSMs) [12], especially Mamba [13], as an emerging sequence model, have become increasingly popular due to their global receptive field and linear complexity. Mamba has demonstrated remarkable performance in many language and vision tasks [14, 15] and is beginning to show promise in remote sensing field [16, 17, 18, 19].
@@
Inspired by the above insight, we propose a novel RSCaMa model to achieve efficient joint spatial-temporal modeling through iterative bi-temporal feature refinement of multiple CaMa layers. Specifically, we introduce Mamba into the RSICC and propose Spatial Difference-aware SSM (SD-SSM) to enhance spatial change perception through differential features. Considering the potential correlation between the temporal scanning characteristics of Mamba and the temporality of the RSICC, to facilitate temporal modeling, we propose the Temporal-Traversing SSM (TT-SSM), which scans bi-temporal features in a temporal cross-wise manner, enhancing the model’s temporal understanding and information interaction. We experimentally confirm the effectiveness of RSCaMa’s efficient joint spatial-temporal modeling and demonstrate the outstanding performance and potential of the Mamba in the RSICC. Additionally, we compare the effect of three language decoders: Mamba, GPT-style decoder with causal attention, and Transformer decoder with cross-attention, providing valuable insight for future RSICC research.
@@
  • To facilitate efficient joint spatial-temporal modeling, we propose RCaMa, which introduces SSM into the RSICC and uses multiple CaMa layers to perform spatial change perception and temporal interaction iteratively.
  • We propose SD-SSM to enhance spatial change perception. To facilitate temporal modeling, we leverage the temporal scanning characteristics of Mamba and propose TT-SSM with a temporal cross-wise scanning manner.
  • The experiment demonstrates RSCaMa’s effective joint spatial-temporal modeling and Mamba’s potential in the RSICC. Besides, our evaluation of language decoding schemes provides valuable insights for future research.
Figure 1: The illustration of the proposed RSCaMa, which consists of three main components: the backbone, multiple CaMa layers, and the language decoder. The CaMa layers play a pivotal role in facilitating efficient joint spatial-temporal modeling. Specifically, SD-SSM focuses on enhancing spatial change perception, while TT-SSM concentrates on temporal interaction.
@@ II-A Preliminaries
$h^{\prime}(t)=\mathbf{A}h(t)+\mathbf{B}x(t),$ $y(t)=\mathbf{C}h(t)$ (1)
where $h^{\prime}(t)$ refers to the time derivative of $h(t)$, ${A}\in\mathbb{R}^{N\times N}$ is the evolution parameter, $\mathbf{B}\in\mathbb{R}^{N\times 1}$ and $\mathbf{C}\in\mathbb{R}^{1\times N}$ serve as the projection parameters.
@@
$\bar{\mathbf{A}}=\mathbf{exp(\Delta A)},$ $\bar{\mathbf{B}}=\mathbf{(\Delta A)}^{-1}\mathbf{(exp(\mathbf{\Delta}A)-I)\cdot\mathbf{\Delta}B}\approx\mathrm{\mathbf{\Delta}}B$ (2)
@@
$h_{k}=\bar{\mathbf{A}}h_{k-1}+\bar{\mathbf{B}}x_{k},$ $y_{k}=\mathbf{C}h_{k}$ (3)
Mamba [13] breaks the LTI limitation of SSM by introducing a selective state space mechanism. It allows the matrices $\bar{\mathbf{B}}$, ${\mathbf{C}}$, and $\mathbf{\Delta}$ to vary with the input sequences, enhancing selective information processing across sequences, enabling the model to be contextually aware of the input.
@@ II-B RSCaMa
As depicted in Fig. 1, we present the overall architecture of the proposed RSCaMa. RSCaMa consists of three main components: the backbone, multiple CaMa layers, and the language decoder. Following prior research [10], we employ the CLIP image encoder that has been image-text aligned as the backbone to extract bi-temporal features, which are then tokenized as $I_{T_{i}}\in\mathbb{R}^{L\times D}(i=1,2)$. $L$ and $D$ are the number and dimensions of tokens respectively. Subsequently, they are added with position embeddings and fed into multiple CaMa layers. The CaMa layers leverage the synergistic effect of SD-SSM and TT-SSM to facilitate efficient joint spatial-temporal modeling, with SD-SSM enhancing spatial change perception and TT-SSM concentrating on temporal interaction. The two token sequences undergo iterative spatial and temporal processing by multiple CaMa layers. Then they are input into a projection layer, bridging the gap between the visual and language domains. Finally, the language decoder utilizes the resulting embeddings to generate descriptive sentences. For the RSCaMa with $N$ CaMa layers, the flow of caption generation is expressed as follows:
${I_{T_{1}}^{(0)},I_{T_{2}}^{(0)}}$ $=\mathrm{\Phi_{pos}}(I_{T_{1}},I_{T_{2}})$ (4)
@@
$\mathrm{T_{cap}}$ $=\mathrm{\Phi_{Dec}}(I_{F})$ (7)
where [;] denotes feature concatenation operation on the channel dimension. $\mathrm{\Phi_{pos}}(\cdot)$ represents adding learnable positional embeddings. $\mathrm{\Phi_{Proj}}(\cdot)$ denotes a projection layer consisting of a linear layer and a residual convolution block.
@@ II-C SD-SSM
Fig. 2 presents the structure of our proposed SD-SSM. The standard Mamba is designed for one-dimensional sequences. Its causal scanning mechanism makes it unsuitable for spatial-aware understanding in visual tasks. Inspired by [20], we perform Forward-SSM and Backward-SSM scans on flattened visual tokens to achieve bidirectional modeling, enhancing understanding of visual features. Furthermore, to improve the model’s spatial perception of changes, we multiply the bi-temporal differencing features and the output of bidirectional SSMs to guide the model. Let’s denote bi-temporal features input to SD-SSM as $I_{T_{1}}$ and $I_{T_{2}}$, the processing of SD-SSM can be represented as follows:
$I^{o}_{T_{i}}$ $=\mathrm{\Phi_{L}}(I^{f}_{T_{i}}*I_{d}+I^{b}_{T_{i}}*I_{d})+I_{T_{i}},i=1,2$ (8)
@@
Figure 3: The structure of the TT-SSM, which rearranges two sequences in a bi-temporal token-wise interleaving manner to facilitate temporal modeling.
@@ II-D TT-SSM
SD-SSM enhances the model’s spatial change perception ability to capture changes sharply using differential features. Furthermore, we propose TT-SSM to facilitate temporal modeling. Considering the potential correlation between the temporal scanning characteristics of Mamba and the temporality of the RSICC, we propose a novel temporal cross-wise manner to scan bi-temporal token sequences. Fig. 3 shows the structure of the TT-SSM. Specifically, for the two sequences $I_{T_{1}}\in\mathbb{R}^{L\times D}$ and $I_{T_{2}}\in\mathbb{R}^{L\times D}$ from SD-SSM, we rearrange them in a bi-temporal token-wise interleaving manner to construct sequences $I_{s}=[I_{T_{1}}^{(0)},I_{T_{2}}^{(0)},I_{T_{1}}^{(1)},I_{T_{2}}^{(1)},I_{T_{1}}^{(2)},...]\in\mathbb{R}^{2L\times D}$ that effectively facilitate temporal information exchange. Finally, $I_{s}$ will be processed by bidirectional SSM and subsequently split into bi-temporal sequences in the original order. This elegant strategy facilitates bi-temporal information interaction. After that, the bi-temporal sequences will be sent to the next CaMa layer, which iteratively performs spatial and temporal processing.
@@ III-A1 Dataset
LEVIR-CC [5] is a large-scale RSICC dataset comprising 10,077 pairs of images and 50,385 descriptive captions. These images capture changes over time in 20 distinct regions across Texas, USA, with a temporal range spanning from 5 to 15 years.
@@ III-A2 Evaluation Metrics
We employ four quantitative metrics to measure the captioning accuracy [21], including BLEU-N (N=1,2,3,4), ROUGEL, METEOR, and CIDEr-D. Higher metric scores indicate better model performance. Besides, following [10], we adopt an average metric $S^{*}_{m}$ as follows:
@@ III-B2 Implementation Details
We implemented all models using the PyTorch deep learning framework and trained them on the NVIDIA GTX 4090 GPU. During training, we use the Adam optimizer to optimize the model parameters. The dimension of the word vector is 768. The number of CaMa layers is set to 3. Our final RSCaMa uses Transformer decoder as the language decoder because of better performance.
@@ III-C Comparison to State-of-the-Art
Currently, mainstream methods commonly utilize CNNs and Transformers as basic design components of the “neck” to process bi-temporal features extracted by the backbone. Methods like Capt-Rep-Diff [6], Capt-Att [6], Capt-Dual-Att [6], and DUDA [6] combine CNNs and attention mechanisms to process bi-temporal features and employ LSTM for language decoding. MCCFormer-S [7], MCCFormer-D [7], RSICCFormer [5], PSNet [9] and PromptCC [10] use Transformer as the core component to construct the neck and employ Transformer’s decoder to generate captions. In this section, we compare RSCaMa with these state-of-the-art (SOTA) methods on model performance, number of parameters, and floating-point operations (FLOPs).
As shown in Table I, RSCaMa demonstrates outstanding performance, particularly excelling in key metrics such as BLEU-4 and $S^{*}_{m}$. Specifically, compared to the latest PromptCC, our model shows +1.70% on BLEU-4 and +1.11% on $S^{*}_{m}$. This significant improvement is attributed to our innovative spatial-temporal modeling units: SD-SSM and TT-SSM. RSCaMa exploits their collaborative effect to promote efficient joint spatial-temporal modeling, which is different from previous methods that use CNN and Transformer to focus on spatial-aware modeling. Besides, our RSCaMa maintains leading performance while having a low number of parameters and computational complexity. Experimental results also suggest that Mamba, as an emerging deep-learning model, has the potential to become a significant option in future model designs. Further exploration of combinations of Mamba with CNNs, Transformers, and Multilayer Perceptrons (MLPs) will be promising research to advance the RSICC technology.
TABLE I: Comparisons experiments on model performance, number of parameters, and floating-point operations (FLOPs). Unlike previous methods relying on CNNs or Transformers, and focusing on spatial-aware modeling, our RSCaMa uses SD-SSM and TT-SSM to facilitate efficient joint spatial-temporal modeling.
Type Method BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGEL CIDEr-D $S^{*}_{m}$ Param. FLOPs
CNN based Capt-Rep-Diff[6] 72.90 61.98 53.62 47.41 34.47 65.64 110.57 64.52 73.21M 19.82G
Capt-Att[6] 77.64 67.40 59.24 53.15 36.58 69.73 121.22 70.17 73.60M 19.90G
Capt-Dual-Att[6] 79.51 70.57 63.23 57.46 36.56 70.69 124.42 72.28 75.58M 20.03G
DUDA [6] 81.44 72.22 64.24 57.79 37.15 71.04 124.32 72.58 80.31M 20.28G
Transformer based MCCFormer-S[7] 79.90 70.26 62.68 56.68 36.17 69.46 120.39 70.68 162.55M 25.09G
MCCFormer-D[7] 80.42 70.87 62.86 56.38 37.29 70.32 124.44 72.11 162.55M 25.09G
RSICCFormer[5] 84.72 76.27 68.87 62.77 39.61 74.12 134.12 77.65 172.80M 27.10G
PSNet[9] 83.86 75.13 67.89 62.11 38.80 73.60 132.62 76.78 319.76M 13.78G
PromptCC[10] 83.66 75.73 69.10 63.54 38.82 73.72 136.44 78.13 408.58M 19.88G
SSM based RSCaMa 85.79 77.99 71.04 65.24 39.91 75.24 136.56 79.24 176.90M 13.03G
TABLE II: Ablation studies on the SD-SSM and TT-SSM. TT-SSM* means that two sequences are concatenated on length instead of the proposed arrangement.
Method SD-SSM TT-SSM TT-SSM* BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGEL CIDEr-D $S^{*}_{m}$
@@
85.64 77.77 70.72 64.86 39.44 74.51 134.97 78.45
III-D Ablation Studies
III-D1 CaMa Layers
The CaMa layers consist of SD-SSM and TT-SSM and play a pivotal role in facilitating efficient joint spatial-temporal modeling. In Table II, the structure of the baseline is similar to SD-SSM, but the difference lies in using self-features rather than differential features to multiply with the output of bidirectional SSM. TT-SSM* means we concatenate two sequences on length instead of adopting the proposed token-wise cross-scanning manner. Experimental results validate that SD-SSM and TT-SSM significantly contribute to improving change captioning performance. This is because SD-SSM effectively enhances the model’s understanding of spatial changes by multiplying differential features with the output of bidirectional SSM, while TT-SSM promotes effective temporal interaction of bi-temporal features through a token-wise cross-scanning manner. The comparison between TT-SSM and TT-SSM* demonstrates that the effectiveness of TT-SSM stems from the proposed cross-scanning manner rather than from the addition of extra bidirectional SSM layers.
We also conducted experiments on the number of CaMa layers. As shown in Table III, our method achieves the best performance when we set the number of CaMa layers to 3.
TABLE III: Parametric experiments on the number of CaMa layers.
Num. BLEU-4 METEOR ROUGEL CIDEr-D $S^{*}_{m}$
2 64.78 39.77 74.75 136.16 78.87
3 65.24 39.91 75.24 136.56 79.24
4 64.77 39.79 74.86 136.72 79.04
TABLE IV: Ablation studies on three language decoders.